OpenStudy (anonymous):

HELPP rationalize the denominator √4/√2

OpenStudy (unklerhaukus):

is that \[\frac{\sqrt4}{\sqrt2}\] ? or \[\sqrt\frac{4}{\sqrt2}\]

OpenStudy (anonymous):

first one

OpenStudy (unklerhaukus):

\[\frac{\sqrt4}{\sqrt2}=\frac{\sqrt4}{\sqrt2}\times1\\\qquad =\frac{\sqrt4}{\sqrt2}\times\frac{\sqrt 2}{\sqrt 2}\\\qquad=\]

OpenStudy (anonymous):

is that the answer?? √2/√2

OpenStudy (anonymous):

or √2/2

OpenStudy (unklerhaukus):

not quite, , what do you get when you multiply the denominators ?

OpenStudy (anonymous):

√2

OpenStudy (anonymous):

im trying to find the X intercept when y=0 for y=-2(x+3)^2 + 4

OpenStudy (goformit100):

I @goformit100 an OS User, Welcomes you to Open Study. 'Discover. Learn. Earn a Certificate Ask questions, give help, and get certified with 150,000 students from 180 countries. ALL THE BEST.

OpenStudy (anonymous):

is there any reason it is not easier just to write \[\frac{\sqrt4}{\sqrt2}=\sqrt{\frac{4}{2}}=\sqrt2\]

OpenStudy (anonymous):

thank u :)